OpenStudy (anonymous):

Erica is in training for a cross country race. She has been training by running 3 miles each day. She will be increasing her running by this weekly formula:y=3(1.1)^3 , where x represents the number of weeks and y represents the distance run. What distance, in miles, will Erica be running in 10 weeks? Round to the nearest mile. 7 8 9 10

OpenStudy (anonymous):

where is the x in the formula? just substitute x = 10 to get the distance y

OpenStudy (anonymous):

where would I distribute the x=10

OpenStudy (anonymous):

you typed the formula in the question wrong. is it \[y = 3*1.1^x\]? substitute x = 10 \[y = 3*1.1^{10}\]

OpenStudy (anonymous):

I got 3.33????

OpenStudy (anonymous):

wait never mind Is the answer 7

OpenStudy (anonymous):

it's 8 7.78 rounds to the nearest whole number is 8

OpenStudy (anonymous):

oh okay thats what I did wrong thanks